A chemist makes 850.mLof silver(II) oxide AgOworking solution by adding distilled water to 170.mLof a 153μM stock solution of silver(II) oxide in water.Calculate the concentration of the chemist's working solution. Be sure your answer has the correct number of significant digits.




1
Expert's answer
2021-12-10T03:27:01-0500

M1V1=M2V2

850×X=170×153

850X=26010

X=26010/850

=30.6μM